OpenStudy (anonymous):

Find the perimeter of a square where one side = 7 inches. do we add all the sides?

OpenStudy (phi):

yes

OpenStudy (king):

yeah...a+a+a+a=4a=4*7=28

OpenStudy (phi):

there are 4 = sides